# Propagation speeds of relativistic conformal fluids from a generalized relaxation time approximation

### v1

@@ II Kinetic theory
Thus we get the linearized equation
$p^{\mu}p^{\nu}\beta_{\nu,\mu}\left(1+\chi\right)+p^{\mu}\chi_{,\mu}={I}_{coll}\left[\chi\right]$ (13)
For the Boltzmann equation, the operator $K$ is symmetric in the space of momentum functions with the inner product [61, 62, 63]
@@
It has exactly four null eigenvectors corresponding to the hydrodynamic modes $\chi_{\mu}=p_{\mu}$; this enforces energy momentum conservation. Now the $H$ theorem requires
$\int Dp\;f_{0}\beta_{\mu}p^{\mu}{I}_{coll}\left[\chi\right]+\int Dp\;f_{0}\chi{I}_{coll}\left[\chi\right]\leq 0$ (16)
The first term vanishes by the symmetry of the linearized collision term. This implies that all non zero eigenvalues of the collision operator must be negative.
The RTA is an attempt to retain these features of the linearized Boltzmann equation but with a simpler collision operator. The first relativistic RTA was Marle’s [18, 19], who wrote the collision operator as
@@
We added a power of $T$ so that $\tau$ should have units of time. Energy momentum conservation requires
@@
where $\rho=3T^{4}/\pi^{2}$ is the energy density for a conformal fluid; we see that in the Anderson-Witting formulation $f_{0}$ is the equilibrium solution in the Landau-Lifshitz frame[65]. While the Anderson-Witting equation is an improvement over Marle’s, both seriously distort the Boltzmann dynamics, and are actually disfavoured by experimental data from relativistic heavy ion collisions [22, 23].
III Generalized RTA
It is clearly seen that trying to improve on Marle’s or Anderson-Witting’s equations by allowing $\tau$ to be momentum-dependent, while keeping the Eckart or Landau-Lifshitz prescriptions to identify the inverse temperature vector, leads to a contradiction. Moreover, if an observer had to know the collision term to choose the “right” prescription (Eckart for Marle’s, Landau-Lifshitz for Anderson-Witting’s) then he could not be doing a true hydrodynamic description; the frame prescription must be given once and for all, in purely hydrodynamic terms. In this section we shall review the collision term proposed in [24, 25], which overcomes these difficulties. For simplicity we shall work in the Landau-Lifshitz frame throughout; this a purely hydrodynamic choice, as it can be determined from the properties of the macroscopic energy-momentum tensor alone.
We write the nonequilibrium 1-pdf as in eq. (10), where $f_{0}$ is the Maxwell-Jüttner distribution eq. (3) built from the Landau-Lifshitz temperature and velocity.
@@ III Generalized RTA
Then by definition
@@
We write the collision integral as in eq. (11). Energy momentum conservation requires ${I}_{coll}\left[\chi\right]$ to be ortogonal to the four null eigenvectors $p^{\mu}$. To satisfy this requirement we introduce a projection operator $Q$ such that for any $g$
@@
$T_{0\nu\rho}^{-1}=\frac{1}{\rho}\left[u_{\mu}u_{\nu}+3\Delta_{\mu\nu}\right]$ (28)
@@
Where $\varsigma$ is a dimensionless relaxation time and $F=F\left[-\beta_{\mu}p^{\mu}\right]$ is a dimensionless function. The entropy production to quadratic order in $\chi$
@@
III.1 Spectral considerations
@@
Now assume that $F\left[p^{\mu}\right]=\lambda$ for some $p^{\mu}_{\lambda}$. Let us work in the Landau-Lifshitz rest frame where $u^{\mu}=\left(1,0,0,0\right)$. Then the solution to eq. (37) is
@@
then from $\left\langle p^{\mu}\zeta_{\lambda}\right\rangle=0$ we conclude that the $c_{\mu}$ themselves are zero. Otherwise we get a linear equation from which we determine the $c_{\mu}$ coefficients.
We see that we may easily find a function $F$ to match any preordained spectrum for the collision operator.
III.2 Marle and Anderson-Witting
To conclude this section, we shall discuss whether it is possible to regard the Marle eq. (17) and Anderson-Witting eq. (19) equations as particular cases of the collision term eq. (29)
To make contact with Anderson-Witting’s equation eq. (19) we set $F=F^{\left(AW\right)}\left[x\right]=x$. Recall that since we are defining the velocity and temperature according to the Landau-Lifshitz prescription, when we split the 1-pdf as in eq. (10), eq. (22) follows, and we get $\left\langle p^{\mu}F\chi\right\rangle=-\beta_{\nu}\left\langle p^{\mu}p^{\nu}\chi\right\rangle=0$, so $Q\left[F\chi\right]=F\chi$.
Assume we also have $\left\langle p^{\rho}\chi\right\rangle=0$ besides eq. (22). Then $Q\left[\chi\right]=\chi$, and so $Q\left[FQ\left[\chi\right]\right]=Q\left[F\chi\right]=F\chi$, yielding the Anderson-Witting RTA. In the following we shall refer to $F\left[x\right]=x$ as the Anderson-Witting prescription.
@@ III.2 Marle and Anderson-Witting
IV The second order Chapman-Enskog solution
As we have mentioned in the Introduction, the first (and probably main) challenge in building hydrodynamics from kinetic theory is to find a suitable parameterization of the kinetic 1-pdf. In this paper we shall take the point of view that the parameterized 1-pdf should take the form eqs. (1) and (2), where the $X_{n}$ are the same functions of momentum as they appear in a Chapman-Enskog solution at some given order.
The Chapman-Enskog solution is a systematic expansion of the 1-pdf in powers of the dimensionless relaxation time $\varsigma$ introduced in eq. (29) [43]. We therefore have a hierarchy of solutions, depending on to which order we extend the expansion.
We use a fiducial temperature $T_{0}$ to build explicit non-dimensional quantities, namely, we define $t=T/T_{0}$ and similarly we adimensionalize all other quantities by dividing or multiplying by $T_{0}$ as required. For simplicity we shall not introduce new names for the dimensionless quantities. The dimensionless Boltzmann equation eq. (29) reads
$p^{\mu}\frac{\partial}{x^{\mu}}f=-f_{0}\frac{{Q}\left[t^{2}F\left[-\beta_{\nu}p^{\nu}\right]{Q}\left[\chi\right]\right]}{\varsigma}$ (44)
@@ IV The second order Chapman-Enskog solution
The Chapman-Enskog procedure aims to obtain a solution of this equation as a expansion in powers of $\varsigma$. Space derivatives in the left hand side of eq. (46) are considered to be of “order zero”, while time derivatives, defined as $\dot{X}=u^{\mu}X_{,\mu}$ have their own development in powers of $\varsigma$

### v2

@@ 3 Relaxation time approximation
Linearization of the Boltzmann collision term yields a $K$ operator which is symmetric in the space of momentum functions with the inner product [75, 20, 21]
@@
${I}_{coll}\left[\chi\right]$ has exactly four null eigenvectors corresponding to the hydrodynamic modes $\chi_{\mu}=p_{\mu}$; this enforces energy momentum conservation. Moreover the $H$ theorem eq. (13) requires that all non zero eigenvalues of the collision operator must be negative.
We shall call a kinetic equation with a collision term as in eq. (19) a generalized relaxation time approximation. The first relativistic RTA was Marle’s [14, 15], who wrote the collision operator of the form
@@
We must mention that Marle’s original expression has a mass $m$ and not a temperature $T$. But as we work with massless particles, the only relevant dimensionful parameter is $T$. Energy momentum conservation requires
@@
where $\rho=3T^{4}/\pi^{2}$ is the energy density for a conformal fluid; we see that in the AW formulation $f_{0}$ is the equilibrium solution in the Landau-Lifshitz frame [77]. Both Marle’s and AW choices seriously distort the Boltzmann dynamics, and are actually disfavoured by experimental data from relativistic heavy ion collisions [23, 24].
3.1 Generalized Relaxation time approximation
Concretely, our concern is to go beyond the Marle and Anderson-Witting RTAs by allowing the relaxation time to depend on the energy of the particle in nontrivial ways. It is clearly seen that trying to improve on Marle’s or AW’s equations by allowing $\tau$ to be momentum-dependent, while keeping the Eckart or Landau-Lifshitz prescriptions to identify the inverse temperature vector, leads to a contradiction [25, 26]. In this section we shall review the collision term proposed in [27, 28], which overcomes this difficulty. For simplicity we shall work in the Landau-Lifshitz frame throughout; this has the appealing feature that it may be determined from the properties of the macroscopic energy-momentum tensor alone.
@@ 3.1 Generalized Relaxation time approximation
where $f_{0}$ is the Maxwell-Jüttner distribution eq. (2) built from the Landau-Lifshitz temperature and velocity. We shall assume the constraint
@@
which follows from applying the Landau-Lifshitz prescription to linear order in the deviation from equilibrium $\chi$.
We write the collision integral as in eq. (12). Energy momentum conservation requires ${I}_{coll}\left[\chi\right]$ to be orthogonal to the four null eigenvectors $p^{\mu}$. To satisfy this requirement we introduce a projection operator $Q$ such that for any $g$
@@
$T_{0\mu\nu}^{-1}=\frac{1}{\rho}\left[u_{\mu}u_{\nu}+3\Delta_{\mu\nu}\right]$ (33)
@@
Where $\varsigma$ is a dimensionless relaxation time and $F=F\left[-\beta_{\mu}p^{\mu}\right]$ is a dimensionless function. The entropy production eq. (13) becomes
@@
The rationale of the proposal eq. (34) is to retain the fundamental features of the Boltzmann equation within a single mathematical structure yet keeping it flexible enough to accommodate phenomenological considerations. Among the former, the feature we want to keep is that the linearized Boltzmann operator is a symmetric operator in a certain Hilbert space [75], which may be therefore diagonalized, and whose spectrum bears basic information about the physics of the system, most notably whether we deal with a hard or soft collision term - in the former case the everpresent zero eigenvalue is an isolated eigenvalue, while in the latter it is part of the continuous spectrum [20, 21, 22]. We elaborate on this in the next Subsection 3.2. Note that the double projection operator $Q$ in eq. (34) both makes the collision term symmetric and enforces energy momentum conservation within a single frame regardless of the function $F$. In this paper we choose to work in the Landau frame throughout, and expect to explore different frame choices in forthcoming work [78, 79, 80].
On the other hand, we leave a window open for phenomenology through the choice of the function $F$ in eq. (34). By far the most common choices for $F$ are Marle’s ($F=constant$) and Anderson-Witting’s ($F\propto-u_{\mu}p^{\mu}$) to be discussed in more detail below (Section 3.3). However, phenomenological considerations in the context of RHICs have led to the proposal of more general power laws [23, 24], which in some cases may be systematically derived from the Boltzmann equation [18, 19] and are actively under research [22, 25, 26, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].. In this paper we only consider functions $F$ defined through power laws; even within this restricted class we find propagation speeds are strongly dependent on the function $F$.
3.2 Spectral considerations
@@
[…]
@@
then from $\left\langle p^{\mu}\zeta_{\lambda}\right\rangle=0$ we conclude that the $c_{\mu}$ themselves are zero. Otherwise we get a linear equation from which we determine the $c_{\mu}$ coefficients. Thus we see that we may easily find a function $F$ to match any preordained spectrum for the collision operator.
3.3 Marle and Anderson-Witting
To conclude this section, we shall discuss whether it is possible to regard the Marle eq. (22) and AW eq. (24) as particular cases of the collision term eq. (34)
To make contact with AW’s equation eq. (24) we set $F=F^{\left(AW\right)}\left[x\right]=x$. Recall that since we are defining the velocity and temperature according to the Landau-Lifshitz prescription, when we split the 1-pdf as in eq. (11), eq. (27) follows, and we get $\left\langle p^{\mu}F\chi\right\rangle=-\beta_{\nu}\left\langle p^{\mu}p^{\nu}\chi\right\rangle=0$, so $Q\left[F\chi\right]=F\chi$.
Assume we also have $\left\langle p^{\rho}\chi\right\rangle=0$ besides eq. (27). Then $Q\left[\chi\right]=\chi$, and so $Q\left[FQ\left[\chi\right]\right]=Q\left[F\chi\right]=F\chi$, yielding the AW RTA. In the following we shall refer to $F\left[x\right]=x$ as the Anderson-Witting prescription.
@@ 3.3 Marle and Anderson-Witting
4 Hydrodynamics from the second order Chapman-Enskog solution
As we have mentioned in the Introduction, the first (and probably main) challenge in seeking a parameterized solution to kinetic theory is to find a suitable parameterization of the kinetic 1-pdf. Our proposal is to use the second order Ch-En solution as a template. This means, we shall work out the second order Ch-En solution and then write it as in eq. (14), thus identifying the $X^{\alpha}\left(x,p\right)$ functions. Of course, in the actual solution these functions are multiplied by given coefficients built from $\beta^{\mu}$ and its derivatives. We shall later replace these coefficients by unknown functions $C_{\alpha}$ obeying the equations of motion (17), thus obtaining a parameterization that generalizes the Ch-En solution.
The Ch-En solution is a systematic expansion of the 1-pdf in powers of the dimensionless relaxation time $\varsigma$ introduced in eq. (34) [45]. We therefore have a hierarchy of solutions, depending on to which order we extend the expansion.
In the moments or Grad approximation on the other side, there is no explicit small parameter with respect to which we can do a perturbative expansion. Thus it is not clear a priori how many moments of the 1-pdf must be included to describe a given departure from equilibrium. To circumvent this situation, in this paper we take the point of view that the parameterized 1-pdf should take the form eqs. (14) and (15), where the $X_{n}$ are the same functions of momentum as they appear in a Ch-En solution at some given order.
We use a fiducial temperature $T_{0}$ to build explicit dimensionless quantities, namely, we define $t=T/T_{0}$ and similarly we make all other quantities non-dimensional by dividing or multiplying by $T_{0}$ as required. For simplicity we shall not introduce new names for the dimensionless quantities. The dimensionless Boltzmann equation eq. (34) reads
$p^{\mu}\frac{\partial}{\partial x^{\mu}}f=-f_{0}\frac{{Q}\left[t^{2}F\left[-\beta_{\nu}p^{\nu}\right]{Q}\left[\chi\right]\right]}{\varsigma}$ (49)
@@ 4 Hydrodynamics from the second order Chapman-Enskog solution
where we have linearized the transport term, which is accurate enough for the discussion below.
The Ch-En procedure aims at obtaining a solution of this equation as a expansion in powers of $\varsigma$. Space derivatives in the left hand side of eq. (51) are considered to be of “zeroth order”, while time derivatives, defined as $\dot{X}=u^{\mu}X_{,\mu}$ have their own development in powers of $\varsigma$